OpenStudy (anonymous):

The mass of a substance varies directly with the volume of the substance. The volume of 100 kilograms of the substance is 80 liters. What is the volume, in liters, of 3.2 kilograms of this substance?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (alexandervonhumboldt2):

80/100=x/3.2

OpenStudy (alexandervonhumboldt2):

2.56

OpenStudy (anonymous):

How'd you find the answer?

OpenStudy (alexandervonhumboldt2):

x=80/100*3.2 th proportons are equal as the same substance

OpenStudy (anonymous):

Ohhh ok.